9:45 A – 7:30 P, 9:00 P-11:30 P
- Group meeting: see notes
- send out invites to Banff workshop
- Continue working on writing referee review
- flipped all bottle stocks
- tossed all Mat alpha tub bottles
- currently expanding: Pc/TM1, Ph/Fm7, Scm/Tm3
- downsized: Df/esc,CyO.
- For virgin collection: Pc/TM1 and Scm/Tm3 (to recombine)
- For virgin collection: dl[D]/CyO, da/Gla (to make maternal double hets)
- New stocks: Gla/CyO, Esc/Gla, da/CyO.
- Started new project: ImageShop.fig
- clone the functions of photoshop I use most
- add some simple batch processing (convert a whole group of images from one format to another, downsizing if desired).
- Could add read/extract lsm features
- Added features today
- multiple image load for batch processing
- multiple image load with tabs and image exit buttons.
- Currently on hold until weekend / my time for coding, need to do more science today.
- backup Projects folder to D:/ drive (54 GB)
Embryo Section staining
- RNA hybe failed — no detectable en/inv or AbdA/Ubx/AbdB staining patterns (though substantially more background in 555 and 488)
- RNA FISH protocol removes previous DNA FISH stain
- 5 New slides (restaining previous 3 embryo sections with AATAT + 2 new embryo sections with BXC).